How much tension must a rope withstand if it used to accelerate 1200 kg
car vertically upward at 0.80 m/s2?
1
Expert's answer
2021-09-01T11:20:04-0400

The tension in a rope

"F=m(a+g)\\
=1200\:\rm kg*(0.80+9.8)\:\rm m/s^2=12720\:\rm N"